OpenStudy (anonymous):

######NEED DONE ASAP###### Graph the following equation: -10x + 8y = -64 I need it graphed please

OpenStudy (compassionate):

Hi, -10x + 8y = -64 Add 10x 8y = 10x - 64 Divide by 8 y = (8/10)x - 8 Happy birthday.

OpenStudy (anonymous):

how would i graph that though?

OpenStudy (compassionate):

8/10 is your slope and -8 is your y-intercept.

OpenStudy (anonymous):

k thanks :)